If 10g of a radioactive substance are present initially and 3 years later only 5g remain, how much will be present after 7 years?
Answer by CubeyThePenguin(3113)

10(x)^3 = 5
x^3 = 1/2
x = 0.793

10(0.793)^7 = 1.984 g